OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? 2x2 - 2x - 1 = 0

OpenStudy (campbell_st):

well you would have b = -2, a = 2 and c = -1 so \[b^2 - 4ac = (-2)^2 - 4 \times 2 \times -1\]